# Prospect Theory Fails for LLMs: Revealing Instability of Decision-Making under Epistemic Uncertainty

### v1

@@ Abstract
Prospect Theory (PT) models human decision-making under uncertainty, while epistemic markers (e.g., “maybe”) serve to express uncertainty in language. However, it remains largely unexplored whether Prospect Theory applies to contemporary Large Language Models and whether epistemic markers, which express human uncertainty, affect their decision-making behavior. To address these research gaps, we design a three-stage experiment based on economic questionnaires. We propose a more general and precise evaluation framework to model LLMs’ decision-making behavior under PT, introducing uncertainty through the empirical probability values associated with commonly used epistemic markers in comparable contexts. We then incorporate epistemic markers into the evaluation framework based on their corresponding probability values to examine their influence on LLM decision-making behaviors. Our findings suggest that modeling LLMs’ decision-making with PT is not consistently reliable, particularly when uncertainty is expressed in diverse linguistic forms. Our code is released in https://github.com/HKUST-KnowComp/MarPT.
@@
Rui Wangthanks: Equal contribution.    Qihan Lin    Jiayu Liu    Qing Zong    Tianshi Zheng    Weiqi Wang    Yangqiu Song Hong Kong University of Science and Technology Huazhong University of Science and Technology rwangcn@connect.ust.hk       yqsong@cse.ust.hk
Figure 1: An overview of our three-stage experiment. Stage 1 fits PT parameters from binary choices with precise probabilities. Stage 2 estimates each marker’s internal probability from the point where options K and U are chosen equally. Stage 3 substitutes probabilities with markers to measure the effect of linguistic uncertainty.
@@ 1 Introduction
Human decision-making under uncertainty has been extensively studied through both normative and descriptive frameworks, among which Prospect Theory (PT) Kahneman and Tversky (1979); Tanaka et al. (2010); Rathi et al. (2025) stands out as particularly influential. PT models a range of behavioral patterns by introducing psychologically grounded parameters, such as loss aversion and probability weighting P. and H. (2005); Gneezy et al. (2006); Rathi et al. (2025). Given the increasing deployment of LLMs in uncertainty-sensitive applications Wei et al. (2025, 2024); Phan et al. (2025), it is crucial to examine whether their decision behaviors align with PT, as such alignment has been shown to enhance human-AI collaboration efficiency Willems et al. (2019); Yao et al. (2024).
In parallel, human uncertainty expressions are commonly expressed via epistemic markers to support flexible language use Belem et al. (2024); Liu et al. (2025b); Zhou et al. (2023). Despite their ubiquity, these expressions are inherently ambiguous Agata (2017), and correctly interpreting them is essential for effective decision-making Mahmoodi et al. (2020). Traditional decision-making frameworks Laplace (1812); Bernoulli (1738); Von Neumann and Morgenstern (1944); Roy (1968) often fail to account for this form of linguistic uncertainty, resulting in a mismatch between computational models and real-world reasoning. While humans are capable of intuitively navigating such expressions Wallsten (1986); Erev and Cohen (1990); Kadavath et al. (2022), it remains unclear how LLMs interpret and respond to them in decision-making contexts.
Existing studies Jia et al. (2024); Liu et al. (2025a, 2024) often assume that LLMs’ decision-making can be directly modeled using PT, without systematically evaluating the suitability of this human-centered psychological framework. Moreover, it remains unexplored whether LLMs exhibit consistent decision-making behavior when probabilities are conveyed through epistemic markers, as is common in real-world communication.
@@
To examine how LLMs handle decision-making under uncertainty, we design a three-stage lottery-based evaluation adapted from classic behavioral economics Allais (1953); Brandstätter et al. (2006); Tanaka et al. (2010). Our evaluation consists of three stages, illustrated in Figure 1. We begin by presenting binary choice tasks with precise numerical probabilities to estimate each model’s risk preferences and fit the PT parameters. In the second stage, we introduce linguistic uncertainty by replacing the numerical probabilities with epistemic markers, allowing us to infer the model’s implicit probabilistic interpretation of these expressions. Finally, we reassess model behavior on the original decision tasks, now framed using epistemic markers grounded in their inferred probability values, to directly evaluate the impact of linguistic uncertainty on decision-making.
@@
Our results reveal three key insights. First, smaller models fail to capture PT-like behavior, whereas larger models (e.g., Qwen2.5-32B-Instruct) demonstrate better alignment with PT predictions. Second, models vary substantially in their interpretations of epistemic markers, resulting in divergent probability mappings. Third, introducing epistemic markers disrupts decision consistency and alters PT parameters, exposing the fragility of LLMs’ decision-making under linguistic uncertainty. These findings point to fundamental gaps in the interpretability and robustness of current LLMs when processing ambiguous language, and underscore the need for improved epistemic calibration and more nuanced theoretical frameworks for modeling LLM behavior in uncertain contexts.
@@ 2 Related Work
Recent work evaluates how LLMs handle uncertain scenarios, including economic games, moral dilemmas, and ambiguous instructions Liu et al. (2025a); Jia et al. (2024); Zong et al. (2025); Zheng et al. (2025d). These studies often probe whether LLMs mimic human cognitive biases or align with normative models. We extend this line of inquiry by testing whether linguistic uncertainty conveyed by epistemic markers causes instability of LLM behaviors in decision-making situations.
@@
Recent work investigates how LLMs interpret and respond to linguistic signals related to uncertainty and confidence. Zhou et al. (2023) study how different epistemic markers embedded in prompts affect model predictions. Belem et al. (2024) evaluate LLMs’ interpretation of epistemic markers, finding partial human-like behavior but systematic biases. Liu et al. (2025b) further argue that such markers are often unreliable indicators of internal confidence in LLMs. These findings suggest that models may mimic surface linguistic patterns rather than demonstrate true epistemic reasoning. Our study builds on this line of work by examining how LLMs understand epistemic markers in economic contexts with uncertainty.
@@ 3 Preliminary
  • probability weighting ($\gamma$): Agents often exhibit systematic probability distortion, overweighting small probabilities and underweighting high probabilities.
@@
The value function v(x) quantifies subjective satisfaction from outcomes relative to a reference point (zero in this study). For outcomes $x\geq 0$ (gains) and $x<0$ (losses), we adopt the parameterized form:
@@
The probability weighting function formalizes the transformation of objective probabilities $p$ in subjective decision weights. We adopt the Kahneman and Tversky (1979) probability weighting function:
@@ 4 Decision-Making Behavior Evaluation
Based on the above economic frameworks, we mainly adopt the three-series lottery-choice experiment developed by Tanaka et al. (2010) to get a reliable PT parameter measurement. Series 1 and 2 are designed to elicit risk preference ($\sigma$) and probability weighting ($\gamma$), while series 3 is designed for loss aversion ($\lambda$). The prospect settings are shown in Appendix A, and the prompt design is in Appendix B.
@@ 5 Decision-making Scenarios with Uncertainty
One of the most fundamental yet underexplored challenges in aligning AI with human reasoning lies in how uncertainty is communicated and interpreted. This section studies decision-making when verbal probability expressions—epistemic markers—replace numerical probabilities. Section 5.1 estimates their numerical equivalents via a controlled lottery experiment, and Section 5.2 applies these values in the PT framework to re-measure PT parameters. The prompt design is in Appendix B.
@@ 5.1 Probability Mapping of Epistemic Markers
Epistemic markers are inherently vague and context-sensitive, yet they often substitute for precise numerical probabilities in practice. For LLMs, the ability to interpret epistemic markers consistently and meaningfully is critical if they are used as decision-support tools. However, there is little empirical understanding of how LLMs map epistemic markers to numerical probabilities, and whether this mapping is coherent across different models or aligned with human intuition.
No. Epistemic Marker
Probability Mapping
by Human
@@
To address this issue, we design a controlled lottery experiment in an economic decision-making context. Each trial presents the model with a choice between two options. For option K, there is a fixed probability pp% of winning \$100, while $p$ traverses all values in $probs=[5,15,25,35,45,55,65,75,85,95]$. For option U, there is an unknown possibility of winning \$100 which is defined by the different markers. We select 14 commonly used markers from previous work in Belem et al. (2024), and present them in Table 1.
For each marker, we record the number of times the model selects option K, denoted as $\text{NUM}K$. The key idea is: when $\text{NUM}K$ reaches 50% of the total trials for a given probability $p$, the model considers the two options equally attractive. We define the mapping probability $p_{mapping}$ for that marker as the value of $p$ at which this occurs:
@@
Since $p$ is sampled at discrete points, the exact 50% point may fall between two sampled probabilities. To estimate $p_{\text{mapping}}$, we perform linear interpolation between the two nearest points. Let $n_{0}=0.5\cdot sample\_number$ be the target count corresponding to 50% selection of option K, $(p_{x},cnt_{x})$ be the probability-count pair just below $n_{0}$, and $(p_{y},cnt_{y})$ be the pair just above $n_{0}$. As illustrated in Figure 3, by the slope formula
@@
5.2 Re-measurement of Prospect Theory Parameters
After we have established the probability mapping, we pick a pair of epistemic markers which has the closest normalized probability to the original settings. This is to ensure that the probability setting does not change much, so that the condition is controlled. For example, for a model with “somewhat unlikely” mapped to $p_{1}=32\%$ and “highly likely” mapped to $p_{2}=68\%$, we normalize them using the formula:
@@ 5.2 Re-measurement of Prospect Theory Parameters
Table 2: Baseline PT parameter estimation across LLMs. Human reference values are taken from Tanaka et al. (2010). Most LLMs exhibit human-like risk aversion and probability distortion, but show inconsistent and generally weaker loss aversion. We bold the results where MAE $\leq 0.20$ and $R^{2}\geq 0.10$, indicating that PT provides a reliable fit.
6 Results and Findings
@@
Table 3: Switching probabilities (%) for top 7 epistemic markers across models. Models differ substantially in how they interpret high-certainty markers such as almost certain or very likely, with mapping values varying by over 50 percentage points across models, indicating inconsistent internal representations of certainty.
@@
Table 4: Switching probabilities (%) for bottom 7 epistemic markers across models. Several models compress distinct low-certainty markers into similarly low probabilities, especially in larger models like Qwen2.5-32B-Instruct, suggesting a limited sensitivity to fine-grained uncertainty.
@@ 6.1 Risk Preference under Numerical Probabilities
For decision-making under exact probability, we obtain different result compared with previous works by Jia et al. (2024) and Liu et al. (2025a). The key parameter values we obtained are shown in Table 2, and more detail can be found in Appendix C.
Not all LLMs decision-making choices can be explained by Prospect Theory.
@@ 6.2 Cross-Model Comparison of Marker Mappings
Across the top 7 markers, the mapping probabilities vary significantly between models. For example, almost certain is mapped to over 97% in Qwen2.5-32B-Instruct but only less than 83% in Qwen2.5-7B-Instruct, while somewhat likely ranges from 8.9% (Qwen2.5-7B-Instruct) to 48.9% (Mistral-7B-Instruct-v0.3). This highlights that language models, even within the same family, do not share a consistent internal representation of uncertainty expressions.
@@
6.3 Change of Decision-Making Behavior under Epistemic Markers
@@
Epistemic markers sometimes activate a more PT-like decision-making behavior for LLMs.
This may because LLMs are more sensitive to linguistic uncertainty than numeric probability value, showing a different probability interpretation of epistemic markers as we obtain in stage 2. However, this finding does not mean epistemic markers can guarantee a more PT-like decision-making behavior. It further strengthen our claim that LLMs’ decision-making is not stable and explainable.
@@ 7 Conclusion
This study challenges the direct application of human-centric Prospect Theory to LLMs, especially in linguistically uncertain contexts. Through a novel three-stage framework combining behavioral economics and semantic analysis, we show that LLMs (1) may not always fit into Prospect Theory (2) exhibit widely divergent interpretations to epistemic markers under decision-making context (3) fail to maintain stable and explainable risk preferences under epistemic markers despite scale advantages. We show that PT fails to model LLMs with epistemic uncertainty and present the first mapping between markers and economic decisions. Future work may explore AI-specific decision theories beyond human dicision models and calibration methods for epistemic markers.
@@
We deliberately choose non-reasoning models and ask them to directly make their choice, because nearly all models tend to directly calculate the expected utility and choose the option with higher EU if they are allowed to reason. This contradicts with our purpose, which is to obtain their risk attitude under quick decisions. However, avoiding reasoning models limits the applicability of our results. Meanwhile, we do not find a universal pattern for PT parameter changes under epistemic markers. Another limitation is that the probability mapping for epistemic markers are obtained in a slightly different context with the decision-making evaluation test. Additionally, our tests are all done in financial decision making context, which may limits its versatility.
Acknowledgments
We thank the anonymous reviewers and the area chair for their constructive comments. The authors of this paper were supported by the ITSP Platform Research Project (ITS/189/23FP) from ITC of Hong Kong, SAR, China, and the AoE (AoE/E-601/24-N), the RIF (R6021-20) and the GRF (16205322) from RGC of Hong Kong,SAR, China.
Appendix A Lottery Design

### v2

@@ Abstract
Prospect Theory (PT) models the human decision-making tendency under uncertainty. While recent studies have developed some questionnaires to elicit the PT parameters (features evaluating decision-making tendency) to describe decision behavior of Large Language Models (LLM), many of them did not report the performance (or explanatory power) of PT itself for LLMs. Additionally, although PT has been used in many LLM-related fields, few studies have tried to test its robustness under linguistic uncertainty, especially epistemic markers (e.g. "maybe"). To address these research gaps, we design an experiment workflow. We adopt a classic economic questionnaire and perform parameter estimation with performance metrics (e.g. McFadden $R^{2}$). We further let LLMs make binary choices which reflect the internal probability values of epistemic markers. We then incorporate epistemic markers into the questionnaire to examine the robustness of Prospect Theory parameters. Our findings suggest that modeling LLMs’ decision-making with PT is not consistently reliable, and applying Prospect Theory to LLMs is likely not robust under epistemic uncertainty.  [^1] We will release our code and data upon acceptance.
@@
Rui Wangthanks: Equal contribution.    Qihan Lin    Jiayu Liu    Qing Zong    Tianshi Zheng    Weiqi Wang    Baixuan Xu    Yangqiu Song Hong Kong University of Science and Technology Huazhong University of Science and Technology rwangcn@connect.ust.hk       yqsong@cse.ust.hk
Figure 1: An overview of our three-stage experiment. Stage 1 fits PT parameters from binary choices with precise probabilities. Stage 2 estimates each marker’s internal probability from the point where options K and U are chosen equally. Stage 3 substitutes probabilities with markers to measure the effect of linguistic uncertainty markers.
@@ 1 Introduction
LLM is increasingly used in important decision-making tasks such as healthcare and finance Keith and Stent (2019); Lehman et al. (2022). While human decision-making under uncertainty has been extensively studied through both normative and descriptive frameworks, the decision patterns for LLMs remain underexplored. Among the human psychological model used in LLM field, Prospect Theory (PT) Kahneman and Tversky (1979); Tanaka et al. (2010); Rathi et al. (2025) stands out as particularly influential. It models a range of human behavioral patterns by introducing psychologically grounded parameters, such as loss aversion and probability weighting P. and H. (2005); Gneezy et al. (2006); Rathi et al. (2025). Since LLM has gained more and more attention, Prospect Theory continuous to play an important role in its training, testing, and alignment Cheng et al. (2025); Wang et al. (2025a). However, although Prospect Theory is intensively used for LLMs, few studies report the performance or explanatory power of this human psychological model when applying it to LLM.
In addition, decision-making is a highly safety-critical task Leyli-abadi et al. (2025), so we must ensure the robustness of explanation theory such as PT under uncertainty. Existing studies have examined how personality prompting, sociodemographic embedding, role-playing interventions, or post-training can systematically shift LLM’s performance Jia et al. (2024); Liu et al. (2025a); Wang et al. (2025a), but whether and how epistemic uncertainty markers Lee et al. (2025) shift LLM decision-making behaviour described under Prospect Theory remains underexplored.
@@
Human uncertainty expressions are commonly expressed via epistemic markers to support flexible language use Belem et al. (2024); Liu et al. (2025c); Zhou et al. (2023). Despite their ubiquity, these expressions are inherently ambiguous Agata (2017), and correctly interpreting them is essential for effective decision-making Mahmoodi et al. (2020). To examine how LLMs handle decision-making under uncertainty, we design a lottery-based evaluation adapted from classic behavioral economics Allais (1953); Brandstätter et al. (2006); Tanaka et al. (2010). Our evaluation consists of three stages, illustrated in Figure 1. We begin by presenting binary choice tasks with precise numerical probabilities to estimate each model’s Prospect Theory parameters. In the second stage, we let LLM to make binary choices between a numerical probability value and a probability described by epistemic markers. This allows us to infer the model’s implicit probabilistic interpretation of those uncertainty expressions. Finally, we reassess model behavior on the original decision tasks, now framed using epistemic markers grounded in their inferred probability values, to directly evaluate the impact of epistemic markers on decision-making.
@@
Our results reveal two key insights. First, Prospect Theory does not consistently performs well at explaining LLM decision behaviors, and it exhibits model-wise differences where larger scale models outperform the others. Second, introducing epistemic markers disrupts decision consistency and alters PT parameters, exposing the fragility of LLMs’ decision-making under linguistic uncertainty. These findings point to fundamental gaps in the interpretability and robustness of current LLMs when processing ambiguous language, underscoring the need for improved epistemic calibration and more nuanced theoretical frameworks for modeling LLM behavior in uncertain contexts.
@@ 2 Related Work
Recent work evaluates how LLMs handle uncertain scenarios, including economic games, moral dilemmas, and ambiguous instructions Liu et al. (2025a); Jia et al. (2024); Zong et al. (2025b); Zheng et al. (2025d). These studies often probe whether LLMs mimic human cognitive biases or align with normative models. We extend this line of inquiry by testing whether linguistic uncertainty conveyed by epistemic markers causes instability of LLM behaviors in decision-making situations.
@@
Recent work investigates how LLMs interpret and respond to linguistic signals related to uncertainty and confidence. Zhou et al. (2023) study how different epistemic markers embedded in prompts affect model predictions. Belem et al. (2024) evaluate LLMs’ interpretation of epistemic markers, finding partial human-like behavior but systematic biases. Liu et al. (2025c) further argue that such markers are often unreliable indicators of internal confidence in LLMs. These findings suggest that models may mimic surface linguistic patterns rather than demonstrate true epistemic reasoning. Our study builds on this line of work by examining how LLMs understand epistemic markers in economic contexts with uncertainty.
@@ 3 Preliminary
  • probability weighting ($\gamma$): Agents often exhibit systematic probability distortion.
@@
The value function v(x) quantifies subjective satisfaction from outcomes relative to a reference point (zero in this study). For outcomes $x\geq 0$ (gains) and $x<0$ (losses), the value function is:
@@
The probability weighting function formalizes the transformation of objective probabilities $p$ in subjective decision weights:
@@ 4 Decision-Making Behavior Evaluation
Based on the above economic frameworks, we mainly adopt the three-series lottery-choice experiment developed by Tanaka et al. (2010) to get a reliable PT parameter measurement. Series 1 and 2 are designed to elicit risk preference ($\sigma$) and probability weighting ($\gamma$), while series 3 is designed for loss aversion ($\lambda$). The prospect settings are shown in Appendix B, and the prompt design is in Appendix C.
@@ 5 Decision-making Scenarios with Uncertainty
Real world decisions are often made under vague linguistic epistemic uncertainty rather than precise numerical probability. This urges for experiments on how epistemic markers influence LLM’s decision-making behaviour. We investigate how decision-making is affected when numerical probabilities are replaced by verbal probability expressions, or epistemic markers. Section 5.1 estimates their numerical equivalents via a controlled lottery experiment, and Section 5.2 applies these values in the PT framework to re-measure PT parameters. The prompt design is in Appendix C.
@@ 5.1 Probability Mapping of Epistemic Markers
Epistemic markers are inherently vague and context-sensitive Liu et al. (2025c); Bergqvist (2015), yet they often substitute for precise numerical probabilities in practice. For LLMs, the ability to interpret epistemic markers consistently and meaningfully is critical if they are used as decision-support tools. However, there is little empirical understanding of how LLMs map epistemic markers to numerical probabilities, and whether this mapping is coherent across different models or aligned with human intuition.
No. Epistemic Marker Probability Mapping by Human
@@
To address this issue, we design a controlled lottery experiment in an economic decision-making context. Each trial presents the model with a choice between two options. For option K, there is a fixed probability pp% of winning \$$M$, while $p$ traverses all values in set probsprobs (see Appendix A.1). For option U, there is an unknown possibility of winning \$$M$ which is defined by the different markers. For the usage of epistemic markers, we manually select 14 commonly used markers from previous work in Belem et al. (2024) to ensure that they are suitable for this context (see Table 1).
For each marker, we record the number of times the model selects option K, denoted as $\text{NUM}_{K}$. The key idea is: when $\text{NUM}_{K}$ reaches $p_{0}=50\%$ of the total trials for a given probability $p$, the model considers the two options equally attractive. We define the mapping probability $p_{mapping}$ for that marker as the value of $p$ at which this occurs:
@@
Since $p$ is sampled at discrete points, the exact $p_{0}$ point may fall between two sampled probabilities. To estimate $p_{\text{mapping}}$, we perform linear interpolation between the two nearest points. Let $n_{0}=p_{0}\cdot sample\_number$ be the target count corresponding to $p_{0}$ selection of option K, $(p_{x},cnt_{x})$ be the probability-count pair just below $n_{0}$, and $(p_{y},cnt_{y})$ be the pair just above $n_{0}$. As illustrated in Figure 3, by the slope formula
@@
gpt-5-mini 0.447 4.000 2.888 0.115 0.206
Gemini-2.5-flash 0.495 1.499 0.814 0.134 0.225
Table 2: Baseline PT parameter estimation across LLMs. Human reference values are taken from Tanaka et al. (2010). For gpt-5-mini and Gemini-2.5-Flash, we use reasoning-enabled prompts while explicitly instructing the models not to compute expected values. We bold the results where MAE $\leq 0.20$ and $R^{2}\geq 0.10$, indicating that PT provides a reliable fit.
@@
gpt-5-mini 97.50 70.81 67.78 50.13 48.12 44.55 04.24
Gemini-2.5-Flash 97.16 61.32 52.34 45.73 32.78 04.94 03.64
Table 3: Switching probabilities (%) for top 7 epistemic markers across models. For gpt-5-mini and Gemini-2.5-Flash, decisions are elicited using reasoning-enabled prompts with an explicit constraint against expected-value computation.
@@
gpt-5-mini 04.16 04.78 02.93 03.17 03.52 02.57 02.58
Gemini-2.5-Flash 03.25 02.86 02.51 02.50 02.52 02.51 02.50
Table 4: Switching probabilities (%) for bottom 7 epistemic markers across models. For gpt-5-mini and Gemini-2.5-Flash, we allow natural reasoning while explicitly prohibiting expected-value calculation.
5.2 Re-measurement of Prospect Theory Parameters
After we have established the probability mapping, we pick a pair of epistemic markers which has a close normalized probability to the original settings. This is to ensure that the probability setting does not change much. Detailed replacement rules are in Appendix D. For example, for a model with “somewhat unlikely” mapped to $p_{1}=32\%$ and “highly likely” mapped to $p_{2}=68\%$, we normalize them using the formula:
6 Results and Findings
@@ 6.1 Risk Preference under Numerical Probabilities
For decision-making under exact probability, we obtain different PT parameters compared with previous works by Jia et al. (2024) and Liu et al. (2025a). This is likely because of using different models and prompt design. The key parameter values we obtained are shown in Table 2, and more detail can be found in Appendix E.
Not all LLMs decision-making choices can be explained well by Prospect Theory.
@@ 6.2 Cross-Model Comparison of Marker Mappings
Across the top 7 markers, the mapping probabilities vary significantly between models. For example, "almost certain" is mapped to over 97% in Qwen2.5-32B-Instruct but only less than 83% in Qwen2.5-7B-Instruct, while "somewhat likely" ranges from 8.9% (Qwen2.5-7B-Instruct) to 48.9% (Mistral-7B-Instruct-v0.3). This highlights that language models, even within the same family, do not share a consistent internal representation of uncertainty expressions.
@@
6.3 Changes of Decision-Making Behavior under Epistemic Markers
@@
Epistemic markers sometimes activate a more human-aligned decision-making behavior for LLMs.
This may because LLMs are more sensitive to linguistic uncertainty than numeric probability value, showing a different probability interpretation of epistemic markers as we obtain in stage 2. However, this finding does not mean epistemic markers can guarantee a more human-aligned decision-making behavior. It further strengthen our claim that LLMs’ decision-making is not stable and explainable.
@@ 7 Conclusion
This study challenges the direct application of human-centric Prospect Theory to LLMs, especially in linguistically uncertain contexts. Through a novel three-stage framework combining behavioral economics and semantic analysis, we show that LLMs (1) decisions may not always explained by Prospect Theory (2) exhibit widely divergent interpretations to epistemic markers under decision-making context (3) fail to maintain stable and reasonable PT parameters under epistemic markers despite scale advantages. We also present the first mapping between markers and probability values in the context of economic decisions.
@@
For the open-source models, we deliberately disable reasoning and require direct choices, as many models otherwise compute expected utilities and deterministically select the option with higher expected value. This design allows us to better isolate risk attitudes under quick decisions. For the closed-source models, we additionally conduct reasoning-enabled experiments with explicit constraints against expected-value computation. Nevertheless, the choice of reasoning settings may still limit the applicability of our results to fully deliberative decision scenarios. Another limitation is that the probability mapping for epistemic markers are obtained in a slightly different context with the decision-making evaluation test.
We do not observe a universal or theoretically interpretable pattern in PT parameter changes under epistemic markers, and thus frame this work as diagnostic rather than mechanistic. Explaining the causes of such instability, exploring more expressive decision theories beyond PT, and developing improved calibration methods for epistemic markers are left for future work.
Ethics Statement
This paper utilize a lottery-based economic questionnaire developed by Tanaka et al. (2010) published on America Economic Association, which allows usage with appropriate citations. The questionaire is done by LLMs, so there is no privacy issues. Our experiment discuss the risk attitude measured under Prospect Theory, which does not contain offensive expressions. The questionnaire is intended to test risk attitude measured by Prospect Theory, and it is used as intended in our paper.
Our experiment involves the usage of Qwen2.5 series models (7B, 14B, 32B) Qwen et al. (2025) with Apache 2.0 license, Mistral-7B-Instruct-v0.3 with Apache 2.0 license, and Llama3.1-8B-Instruct with Llama3.1 license. They run on an 8x RTX 3090 GPU cluster.
Our paper mainly tests the robustness of PT under epistemic uncertainty, which points out risk of using PT in LLM-related fields. This does not introduce extra risks. Our research focuses on financial decision-making within the English language domain.
Appendix A Hyperparameters
A.1 Probability Mapping Experiment
In the probability mapping experiments in 5.1, the monetary reward is fixed at $M=100$. This value is chosen to provide a clear and intuitive payoff magnitude without introducing excessive numerical complexity.
The probability parameter $p$ takes values from the set $probs=\{5,15,25,35,45,55,65,75,85,95\}$. These values are selected to uniformly cover the range of possible probabilities from low to high in increments of 10 percentage points, enabling systematic analysis of internal probability values of epistemic markers.
A.2 Model Generation
Temperature is set to $0.7$ for all experiments, following prior work on decision-making behavior of LLMs under uncertain contexts (Jia et al., 2024; Liu et al., 2025b, 2026). Since our evaluation relies on sampling-based decoding to capture distributional decision behavior, the temperature cannot be set to zero. The chosen value balances diversity and coherence and is commonly adopted in LLM evaluations. All other decoding and generation hyperparameters use the default settings provided by the HuggingFace implementation.
Hyperparameter Value
Generation method Sampling
Temperature 0.7
Maximum new tokens 8
Padding token EOS token
EOS token EOS token
Batch size 16
History length 10
Number of lottery rounds 35
Table 5: Key hyperparameters for model generation
Appendix B Lottery Design